Consider the electric field 𝐸 = π‘Žπ‘Ÿ

π‘ŸπœŒπ‘™

4πœ‹πœ–0

∫

𝑑𝑧

β€²

(π‘Ÿ

2+𝑧′

2)

3/2

+∞

βˆ’βˆž

due to the uniform line charge density ρl

= 12 Β΅C/m of an infinitely long straight conducting wire at a distance r. Find out the numeric value ofΒ 

r. Remember that prime variables corresponds to the source quantity.


1
Expert's answer
2021-05-03T10:28:35-0400

Given,

Electric field "E=ar"

"E=\frac{r\rho l}{4\pi \epsilon_o}\int_{-\infty}^{\infty} \frac{dz}{(r^2+z^2)^{3/2}}"


"=\frac{ar\rho l}{4\pi \epsilon_o}\times [\frac{z}{r^2(z^2+r^2)}]_{-\infty}^{\infty}"


"=\frac{ar\rho l}{4\pi \epsilon_o}\times (\frac{2}{r^2})"

So,

"ar=\frac{r\rho l}{4\pi \epsilon_o}\times (\frac{2}{r^2})"


"\Rightarrow" "r=\sqrt{\frac{2\rho l}{4\pi \epsilon_o a}}"